### v1

# Channel-wise Parallelizable Spiking Neuron with Multiplication-free Dynamics and Large Temporal Receptive Fields

Abstract
Spiking Neural Networks (SNNs) are distinguished from Artificial Neural Networks (ANNs) for their sophisticated neuronal dynamics and sparse binary activations (spikes) inspired by the biological neural system. Traditional neuron models use iterative step-by-step dynamics, resulting in serial computation and slow training speed of SNNs. Recently, parallelizable spiking neuron models have been proposed to fully utilize the massive parallel computing ability of graphics processing units to accelerate the training of SNNs. However, existing parallelizable spiking neuron models involve dense floating operations and can only achieve high long-term dependencies learning ability with a large order at the cost of huge computational and memory costs. To solve the dilemma of performance and costs, we propose the mul-free channel-wise Parallel Spiking Neuron, which is hardware-friendly and suitable for SNNs’ resource-restricted application scenarios. The proposed neuron imports the channel-wise convolution to enhance the learning ability, induces the sawtooth dilations to reduce the neuron order, and employs the bit shift operation to avoid multiplications. The algorithm for design and implementation of acceleration methods is discussed meticulously. Our methods are validated in neuromorphic Spiking Heidelberg Digits voices, sequential CIFAR images, and neuromorphic DVS-Lip vision datasets, achieving the best accuracy among SNNs. Training speed results demonstrate the effectiveness of our acceleration methods, providing a practical reference for future research.
111$\dagger$ Corresponding authors
1 Introduction
Inspired by the biological neural system, Spiking Neural Networks (SNNs) are regarded as the third generation neural network models (Maass, 1997). By emulating neuronal dynamics and spike-based communication characteristics in the brain (Tavanaei et al., 2019), SNNs effectively capture temporal information and achieve event-driven efficient computation, providing a novel paradigm for building the spike-based machine intelligence (Yao et al., 2024b).
Spiking neurons are the key component that distinguishes SNNs from Artificial Neural Networks (ANNs) (Li et al., 2024a). They integrate input currents from synapses to membrane potentials by complex neuronal dynamics and fire spikes when the membrane potentials cross the threshold. These proceedings are generally described by several discrete-time difference equations (Fang et al., 2021b, 2023a) in a formulation similar to the Recurrent Neural Networks. The discrete threshold-triggered firing mechanism induces the nondifferentiable problem and impedes the application of gradient descent methods. Recently, this problem has been resolved to a considerable degree by the surrogate gradient methods (Wu et al., 2018; Shrestha and Orchard, 2018; Neftci et al., 2019).
Deep SNNs (Fang et al., 2021a; Zhou et al., 2023; Yao et al., 2024a) commonly use stateless synapses, i.e. the weights are shared across time-steps and outputs only depend on the inputs at the same time-step, to extract spatial features. Consequently, dynamic spiking neurons play the critical role in SNNs to extract temporal features and have attracted many research interests. Most of the previous research focuses on increasing the neuron model complexity with learnable parameters (Fang et al., 2021b; Yao et al., 2022) or adaptive dynamics (Huang et al., 2024a), which strengths the model capacity but brings extra computation costs. Another issue is that traditional spiking neuron models run in a serial step-by-step mode and cannot fully utilize the powerful massive parallel computing ability of Graphics Processing Units (GPUs), which leads to slower training speed of SNNs than ANNs. Recently, parallelizable spiking neuron models (Fang et al., 2023b; Yarga and Wood, 2023; Huang et al., 2024c; Chen et al., 2024a) have been proposed that overwhelm traditional serial models in running speed, showing a promising solution to accelerate the training of SNNs.
In previous designs of spiking neurons, the computational cost associated with the neuron dynamics is generally overlooked. For instance, the Complementary Leaky Integrated-and-Fire (CLIF) neuron (Huang et al., 2024a) introduces the computationally expensive sigmoid exponentiation, the Parallel Spiking Neuron (PSN), and the masked PSN (Fang et al., 2023b) rely on the dense floating-point matrix multiplication. Compared to PSN and masked PSN, sliding PSN (Fang et al., 2023b) only requires convolutional operations and demonstrates superior performance in handling time series with variable lengths. However, according to the study by (Fang et al., 2023b), sliding PSN only achieves comparable performance to PSN when using a large convolutional kernel size, denoted as the order of the neuron, which significantly increases computational cost and memory usage.
In this article, we focus on designing a new variant of parallelizable spiking neuron models with hardware-friendly dynamics, low computation cost, and high long-term dependency learning ability. We propose an enhanced neuronal architecture named Multiplication-Free Channel-wise Parallel Spiking Neurons (mul-free channel-wise PSN), whose neuronal dynamics are shown in Figure 1, and validate its performance with state-of-the-art (SOTA) accuracy on temporal datasets. Our contributions are as follows.
  • To enhance the temporal information capturing ability, we derive the sliding PSN by applying the channel-wise separable convolutions. To solve the dilemma of large temporal receptive fields and computational costs, we use the dilated convolution. Compared to sliding PSN, our improvement does not introduce any additional floating point operations (FLOPs).
  • To avoid the costly multiplication operations, we supersede them with bit-shift operations, which are hardware-friendly for resource-restricted neuromorphic chips. The theoretical energy cost and area for hardware implementation with 8-bit integers (INT8) precision under 45nm CMOS is reduced 8$\times$, from 0.2 $pJ$ and 282 $\mu m^{2}$ to 0.024 $pJ$ and 34 $\mu m^{2}$ (You et al., 2024), respectively.
  • We deliberate on the implementations of SNNs with mul-free channel-wise PSN on GPUs for efficient training. We propose an autoselect algorithm to choose the fastest implementations, which is practical for future research about accelerating parallelizable spiking neurons.
  • We achieve SOTA performance on various temporal tasks, validating the superior capability of the proposed methods in learning long-term dependencies.
Figure 1: The neuronal dynamics of the mul-free channel-wise Parallel Spiking Neuron.
2 Related Work
2.1 Hardware-friendly Network Design
To deploy neural network models to edge devices such as mobile phones and Field Programmable Gate Arrays with limited energy, memory, and computational ability, a promising solution is hardware-friendly network design. Various methodologies have been proposed. Network quantization (Gholami et al., 2021) quantizes the original weights and activations to low bits. Typical models in 8-bit integers require 4×4$\times$4 × less memory consumption and achieve up to 4×4$\times$4 × faster computation than models in 32-bit floating-points. Network pruning (Cheng et al., 2024) prunes synapses and neurons to reduce the size of the model. Classic pruning methods can achieve 4×4$\times$4 × compression ratio for ResNet-18 on ImageNet with about $5\%$ accuracy drop (Blalock et al., 2020). Knowledge distillation (Gou et al., 2021) employs a large teacher network to supervise the learning of a small student network, and the student network can obtain a competitive or even higher performance than the teacher network. Apart from the above universal methods, SNNs (Maass, 1997; Roy et al., 2019) achieve extreme power efficiency by asynchronous event-driven computation in tailored neuromorphic chips. For instance, Intel Loihi (Davies et al., 2018) consumes 48×48$\times$48 × energy efficiency than CPUs in solving the LASSO optimization problem; Tsinghua Tianjic (Pei et al., 2019) achieves up to $10^{4}$ times power efficiency over the Titan-Xp GPU when classifying the NMNIST (Orchard et al., 2015).
2.2 Spiking Deep Learning
The performance of SNNs surges by the induction of deep learning methods. ANN to SNN conversion (Cao et al., 2015; Hu et al., 2023) and surrogate gradient learning (Neftci et al., 2019) are two primary methods for spiking deep learning (Chollet, 2017). The conversion method leverages the average firing rate of spiking neurons to approximate the continuous activations in ANNs, transforming a pre-trained ANN into an SNN through techniques including threshold adjustment and weight normalization. However, the conversion method requires a substantial number of time-steps to estimate accurate firing rates. The surrogate gradient learning method employs the derivative of a continuous function to replace the derivative of the Heaviside function used in the spike generation process. This approach enables SNNs to achieve credit assignment through backpropagation through time (BPTT) and gradient descent. Compared with the conversion method, the surrogate gradient method requires much fewer time-steps, but the training cost is higher due to the use of BPTT.
2.3 Spiking Neuron Models
The improvement of spiking neuron models provides a general method to upgrade SNNs, which attracts many interests in the research community. The Parametric Leaky Integrated-and-Fire (PLIF) spiking neuron (Fang et al., 2021b) parameterizes the membrane time constant $\tau_{m}$ by a sigmoid function and can learn $\tau_{m}$ by gradient descent during training, showing better accuracy and lower latency than the traditional Leaky Integrated-and-Fire (LIF) neuron with fixed $\tau_{m}$. The Gated LIF (GLIF) neuron (Yao et al., 2022) assembles the learnable gate units to fuse different bio-features in the neuronal behaviors of membrane leakage, integration accumulation, and reset, achieving impressive performance by these rich neuronal patterns. The Complementary LIF (CLIF) neuron (Huang et al., 2024b) introduces the complementary membrane potential into the LIF neuron, effectively capturing and maintaining information related to membrane potential decay. However, the sigmoid used in its neuronal dynamics cannot be removed during inference, which is costly for neuromorphic chips.
The Parallel Spiking Neuron (PSN) family (Fang et al., 2023b) and the Stochastic Parallelizable Spiking Neuron (Stochastic PSN) (Yarga and Wood, 2023) are the first parallelizable spiking neuron models. These two models convert the iterative neuronal dynamics to a non-iterative formulation by removing the neuronal reset. Extending from PSN, several variants are proposed. The Parallel Multi-compartment Spiking Neuron (PMSN) (Chen et al., 2024b) introduces multiple interacting substructures to enhance the learning ability over diverse timescales. The Parallel Spiking Unit (PSU) (Li et al., 2024b) adds a fully-connected layer with sigmoid activations inside the neuron to approximate the neuronal reset. These methods obtain performance gains over PSN in certain datasets, but increase the complexity of neuron models and slow down training speeds.
3 Preliminary
3.1 Vanilla Spiking Neuron
The general formulations of spiking neurons are as follows:
$H[t]$ $=f(V[t-1],X[t]),$ (1)
$S[t]$ $=\Theta\left(H[t]-V_{th}\right),$ (2)
$V[t]$ =$=\begin{cases}H[t]\cdot(1-S[t])+V_{reset}\cdot S[t],\text{hard reset}\\ H[t]-V_{th}\cdot S[t],\text{soft reset}\end{cases}.$ (3)
Eq.(1) illustrates the neuronal charging process, where $X[t]$ is the input current, $H[t],V[t]$ are the membrane potential before charging and after resetting, and $f$ is the charging equation specified for different spiking neurons. After charging, the membrane potential $H[t]$ will be compared with the threshold $V_{th}$, described as Eq.(2) to determine whether firing. $\Theta(x)$ is the Heaviside step function, defined as $\Theta(x)=1$ for $x\geq 0$ and $\Theta(x)=0$ for $x<0$. If firing, the membrane potential will be reset as in Eq.(3). There are mainly two types of reset methods: hard reset will force the $V[t]$ to $V_{reset}$, while soft reset will subtract $V_{th}$ from $V[t]$.
3.2 Parallel Spiking Neuron
Fang et al. (2023b) found that for commonly used spiking neurons with a linear sub-threshold dynamic Eq.(1), such as the Integrate-and-Fire (IF) neuron and the LIF neuron, the neuronal dynamics could be expressed in a non-iterative form after removing the reset equation Eq.(3):
$H[t]=\sum_{i=0}^{T-1}W[t][i]\cdot X[i],$ (4)
where $W[t][i]$ is determined by Eq.(1). For example, $W[t][i]={\tau_{m}}^{-1}(1-{\tau_{m}}^{-1})^{t-i}\cdot\Theta(t-i)$ for the LIF neuron whose neuronal charging equation is:
$H[t]=(1-{\tau_{m}}^{-1})\cdot V[t-1]+{\tau_{m}}^{-1}\cdot X[t].$ (5)
Fang et al. (2023b) extended Eq.(4) by setting $W[t][i]$ as a learnable parameter, and proposed the PSN with following neuronal dynamics:
$\bm{H}$ $=\bm{W}\bm{X},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\bm{W}\in\mathbb{R}^{T\times T},\bm{X}\in\mathbb{R}^{T},$ (6)
$\bm{S}$ $=\Theta(\bm{H}-\bm{V}_{th}),~{}~{}~{}~{}~{}\bm{V}_{th}\in\mathbb{R}^{T},$ (7)
where $T$ is the sequence length. For simplicity, we ignore the batch dimension. The PSN does not involve iteration over time-steps. The core computation of the PSN is the matrix multiplication, which is highly optimized on GPUs and can be computed in parallel. Modified from the PSN, the sliding PSN is proposed by (Fang et al., 2023b) with hidden states generating from the last $k$ inputs by a shared weight $\bm{W}\in\mathbb{R}^{k}$ across time-steps, whose neuronal dynamics are as follows:
$H[t]$ $=\sum_{i=0}^{k-1}W[i]\cdot X[t-k+1+i],$ (8)
$S[t]$ $=\Theta(H[t]-V_{th}),$ (9)
where $X[j]=0$ for any $j<0$ and $k$ is the order of the neuron. The sliding PSN can process input sequences with variable lengths, and the number of its parameters is decoupled with $T$. It can output $H[t]$ at the time-step $t$, while the PSN can only generate outputs after receiving the whole input sequence, making it more suitable for temporal tasks.
4 Methods
4.1 Channel-wise and Dilated Convolution
In PSN and sliding PSN, the weights of the neurons are shared across all channels. However, the visualization of feature maps from an SNN conducted by (Fang et al., 2021b) implies that the difference between channels is huge, e.g., one channel extracts the edges and another channel extracts the background at all time-steps (refer to Figure S4 in (Fang et al., 2021b) for more details). This coarse design concept may fail to capture the subtle disparity of features in channels. To solve this issue, we extend the weights to channel-wise.
To capture long-term dependency, the sliding PSN must use a large order $k$, resulting in a significant rise in computational cost and memory usage. We overcome this issue through the dilated convolution (Yu and Koltun, 2016), where the convolution no longer processes consecutive inputs $(...,X[t-2],X[t-1],X[t])$, but instead $(...,X[t-2d],X[t-d],X[t])$, with $d>1$ as the dilation rate.
Specifically, denote the input sequence as $\bm{X}\in\mathbb{R}^{T\times C}$, where $T$ is the sequence length and $C$ is the number of channels. We propose the channel-wise PSN with the following neuronal charging equation:
$H[t][c]$ $=\sum_{i=0}^{k-1}W[c][i]\cdot X[t-(k-1-i)\cdot d][c],$ (10)
where $\bm{W}\in\mathbb{R}^{C\times k}$ is the learnable weight, $k$ is the order of the neuron and $d\in\mathbb{N}^{+}$ is the dilation rate. The channel-wise PSN has identical FLOPs to the sliding PSN, and the latter can be regarded as a simplified case with setting $W[0][i]=W[1][i]=...=W[C-1][i]$ and $d=1$ in Eq.(10).
Figure 2: The temporal receptive field increases with depths at a slow rate in the sliding PSN with (a) fixed dilations and a fast rate in the channel-wise PSN with (b) sawtooth dilations.
Additionally, when using multiple layers of dilated convolutions, setting the same dilation rate will lead to the grid effect. An approach to solving this issue is to assign the dilation rate using a sawtooth wave-like heuristic (Wang et al., 2018). Specifically, when constructing SNNs with channel-wise PSNs, we start by initializing with $d^{0}=1$ for the first spiking neuron layer, where the superscript represents the spiking neuron layer index. Then we update $d^{l}$ as:
$d^{l+1}=d^{l}\mod 3+1.$ (11)
This approach ensures that after stacking multiple layers, the convolution in the time domain could effectively incorporates inputs from all time-steps. Figure 2 shows how the temporal receptive field increases with depths with (a) fixed dilations $d=1$ in the sliding PSN and (b) sawtooth dilations in the channel-wise PSN. The order is $k=2$ in both cases. It can be found that, with increasing depths, both neurons achieve larger temporal receptive fields. However, the sliding PSN can only capture the last 4 time-steps with 4 layers, while the channel-wise PSN can capture the last 7 time-steps.
4.2 Multiplication-Free Neuronal Dynamics
One of the most attractive characteristics of SNNs is that the multiply-accumulate (MAC) operations between binary spikes and synaptic weights can be superseded by accumulate (AC) operations during inference in neuromorphic chips (Pei et al., 2019). However, computational costs of the neuronal dynamics have not been paid much attention. For example, the PSN, the sliding PSN, and the Stochastic PSN involve massive MAC operations between floating-point neuronal weights and input currents. To future reduce the internal computation costs of spiking neurons, we introduce the bit-shift operation to supersede multiplication, which has been successfully employed in quantized neural networks (You et al., 2020; Elhoushi et al., 2021).
Figure 3: The workflow of power-of-2 quantizer.
When $w$ is the power of 2, then the multiplication of $w$ and $x$ can be replaced by the bit shift operation:
$w\cdot x=x<<\log_{2}(w),$ (12)
where <<much-less-than$<$ < is the left bit shift operation. In particular, when $\log_{2}(w)<0$, left shifting an negative number $\log_{2}(w)$ of bits is actually right shifting $|\log_{2}(w)|$. To employ the bit-shift operation, we quantize $\bm{W}$ in Eq.(10) to $\bm{W}_{q}$, whose elements are the power of $2$, by an quantizing function $\mathcal{Q}$:
$\bm{W}_{q}=\mathcal{Q}(\bm{W})=\mathrm{sign}(\bm{W})\cdot 2^{\mathrm{round}(\log_{2}(|\bm{W}|))},$ (13)
where $\mathrm{sign}(x)$ is the sign function and returns the sign ($1$ or $-1$) of the input $x$; $\mathrm{round}(x)$ is the rounding-to-nearest function. Figure 3 shows the workflow of Eq.(13).
Figure 4: Using the Straight-Through Estimator to redefine $\mathrm{round}^{\prime}(x)$ solely will still cause a discrete gradient $\mathcal{Q}^{\prime}(x)$.
Remarkably, the gradient of $\mathrm{round}(x)$ is zero almost everywhere, and other operations in Eq.(13) are differentiable. The standard practice is to employ the Straight-Through Estimator (Bengio et al., 2013) to redefine its gradient as 1:
$\mathrm{round}^{\prime}(x)=1.$ (14)
Then the gradient of Eq.(13) is:
$\mathcal{Q}^{\prime}(x)=\frac{1}{|x|}\cdot 2^{\mathrm{round}(\log_{2}(|x|))}.$ (15)
However, Eq.(15) is still unstable because $\mathrm{round}(x)$ causes jump points and it oscillates around 0, shown in Figure 4.
To avoid the numerical instability caused by Eq.(15), we redefine $\mathcal{Q}^{\prime}(x)$ as a whole Straight-Through Estimator, rather than using Eq.(14) solely:
$\frac{\partial\bm{W}_{q}}{\partial\bm{W}}=\mathrm{\mathcal{Q}}^{\prime}(\bm{W})=\bm{1}.$ (16)
The neuron model is called mul-free channel-wise PSN when using $\bm{W}_{q}$ in Eq.(10), and the complete neuronal dynamics is as follows, which is illustrated in Figure 1:
$H[t][c]$ $=\sum_{i=0}^{k-1}X[t-(k-1-i)\cdot d][c]<<\log_{2}(W_{q}[c][i]),$ (17)
$S[t][c]$ $=\Theta(H[t][c]-V_{th}[c]),$ (18)
where $\bm{V}_{th}\in\mathbb{R}^{C}$ is the learnable channel-wise threshold, and $\log_{2}(\bm{W}_{q})\in\mathbb{Z}^{C\times k}$ is quantized from $\bm{W}$ by Eq.(13). Note that $\log_{2}(\bm{W}_{q})$ is solved beforehand and there are no $\log$ and multiplication operations during inference.
4.3 Training Acceleration
The motivation for proposing parallelizable spiking neuron models is to solve the slow training speed of SNNs on GPUs caused by the step-by-step iterations of traditional serial neuron models. Thus, the efficient implementation of mul-free channel-wise PSN requires elaborate consideration.
Figure 5: Reshape operations involving adjacent dimensions are free of memory reading/writing and are much faster than those involving nonadjacent dimensions.
The computations of mul-free channel-wise PSN focus on Eq.(17), which is a typical 1-D convolution. The straightforward implementation uses PyTorch’s 1-D convolution (Conv1d) for the proposed neuron layer and the prevailing spiking deep learning framework SpikingJelly (Fang et al., 2023a) for other modules. However, Conv1d requires the sequence length, which is the time dimension, as the last dimension. While the default data layout in SpikingJelly is the time-first layout, using $(T,N,...)$ to represent the sequence data, where $T$ is the length, $N$ is the batch size, and ”...” represents any additional dimensions. This layout is necessary for accelerating stateless layers by fusing time and batch dimensions for parallel computing (Fang et al., 2023a). In this circumstance, the straightforward implementation is:
(1) Time-first + Reshape + Conv1d: this method uses PyTorch’s Conv1d, which requires the time-step at the last dimension. Thus, the reshape operations $(T,N,C,...)\rightleftharpoons(N*,C,T)$ before and after Conv1d are unavoidable, where $N*$ represents any additional dimensions in ”...” have been fused to the $N$ dimension into $N*$.
Unfortunately, the reshape operations in implementation (1) for the proposed neuron involve nonadjacent dimensions, resulting in slow memory reading/writing (R/W). Note that physical memory is 1-D, resulting in the fact that data in nonadjacent dimensions are also nonadjacent in physical memory. This is also the reason why SpikingJelly requires the time-first layout: the time and batch dimension fusion requires reshape operations $(T,N,...)\rightleftharpoons(TN,...)$, which involves adjacent dimensions and gets rid of memory R/W. Figure 5 demonstrates the cases of reshape operations with or without memory R/W.
Different implementations require different memory layouts, and consequentially lead to speed difference, even if the theoretical FLOPs are identical. Thus, the choice of data layouts is the primary principle. Different from the time-first layout in SpikingJelly, the time-last layout uses $(N,...,T)$ to represent sequences, which can also be considered. Based on the aforementioned background, we design the following five more implementations as candidates:
Algorithm 1 Autoselect acceleration algorithm
Require: An SNN stacked with $L$ layers $\big{\{}M_{0},M_{1},...,M_{L-1}\big{\}}$. The layer $M_{l}$ has $n_{l}$ optional acceleration methods. The input sequence $\bm{X}_{0}$.
1:  for $\Omega\leftarrow$time-first, time-last}
2:    Reshape $\bm{X}_{0}$ to $\Omega$
3:    $t_{\Omega}=0$
4:      for $l\leftarrow 0,1,...L-1$
5:        for $i\leftarrow 0,1,...,n_{l}-1$
6:          Record the current time $\mathcal{T}_{0}$
7:          Execute the forward propagation $\bm{Y}_{l}=M_{l}(\bm{X}_{l})$
8:          Record the current time $\mathcal{T}_{1}$
9:          Randomize a tensor $\bm{Z}_{l}$ with the same shape as $\bm{Y}_{l}$
10:          Record the current time $\mathcal{T}_{2}$
11:          Execute the backward propagation $M_{l}^{\prime}(\bm{Z}_{l})$
12:          Record the current time $\mathcal{T}_{3}$
13:          $t_{l,i}$ = $\mathcal{T}_{1}-\mathcal{T}_{0}+\mathcal{T}_{3}-\mathcal{T}_{2}$
14:        Choose the faster method $a_{\Omega,l}=\mathrm{argmin}_{i}(t_{l,i})$
15:        $t_{\Omega}\leftarrow t_{\Omega}+\min(t_{l,i})$
Outputs: The layout $\Omega^{*}=\mathrm{argmin}_{\Omega}(t_{\Omega})$ and the acceleration method $a_{\Omega*,l}$ for $M_{l}$
(2) Time-first + Reshape + Conv2d: this method uses PyTorch’s 2-D convolution (Conv2d). It involves the reshape operations $(T,N,C,...)\rightleftharpoons(N,C,*,T)$ or $(T,N,C,...)\rightleftharpoons(N,C,T,*)$, and sets weight and stride in the ”*” dimension as 1. Compared with the implementation (1), its memory copying cost is less, while the 2-D convolution is more costly.
(3) Time-last + Vmap + Conv1d: this method uses the vectorizing map function (Vmap) in PyTorch to vectorize Conv1d to process the input sequence with the $(N,C,*,T)$ layout over the ”*” dimension. The reshape operations $(N,C,...,T)\rightleftharpoons(N,C,*,T)$ are nearly cost-free because the reshaped dimensions are adjacent physically.
(4) Time-last + Conv2d: this method is similar to the implementation (3), but processes the input sequence with the $(N,C,*,T)$ by Conv2d and sets the weight and stride in the ”*” dimension as 1, rather than by Vmap.
(5) Time-first/last + Vmap + MM: this method uses Vmap to vectorize matrix multiplication (MM) to process inputs over channels ($c$ in Eq.(17)). Refer to the Appendix for more details about how the weights for MM are generated.
(6) Time-first/last + Custom CUDA Kernel: this method avoids reshape operations and can be used for any memory layout. However, the convolutions in PyTorch are highly optimized, i.e. implemented by the official NVIDIA CUDA Deep Neural Network (cuDNN) library, which might be much faster than custom implementations.
Note that the implementations (3)-(6) adopt the time-last layout. Then the stateless layers should also use the same layout. Otherwise, reshape operations between time-first and time-last layouts will cause great latency. Correspondingly, the time batch dimension fusion method to accelerate stateless layers in SpikingJelly cannot be applied. We also elaborate on acceleration methods for stateless layers in the time-last layout, and the details can be found in the Appendix. As the acceleration of mul-free channel-wise PSN on GPUs will influence the stateless layers mutually, and the acceleration effect varies with the input shapes and GPUs, choosing of acceleration methods is empirical. We design an autoselect acceleration algorithm to avoid manually choosing, as shown in Algorithm 1. When the training of an SNN starts, the shape of the input sequence is also determined. Then this algorithm will run a benchmark over layouts and acceleration methods and select the options with the fastest speed.
[…]
Hammouamri et al. Two-layer FC + LIF + Learned Delay 95.10
Li et al. (2024b) Four-layer FC + RPSU 92.49
Chen et al. (2024b) Two-layer FC + PMSN 95.10
Yarga and Wood (2023) Two-layer FC + Stochastic PSN + Learned Delay 95.01
Ours Two-layer FC + Mul-free Channel-wise PSN + Learned Delay 95.71
Table 1: Comparison with the state-of-the-art SNN methods on the SHD dataset.
Datasets Mul-free Channel-wise PSN PMSN PSN masked PSN sliding PSN GLIF PLIF LIF
Sequential CIFAR10 91.17 90.97 88.45 85.81 86.70 83.66 83.49 81.50
Sequential CIFAR100 66.21 66.08 62.21 60.69 62.11 58.92 57.55 53.33
Table 2: Comparison of test accuracy (%) of spiking neurons on sequential CIFAR datasets.
Method Frontend Backend Accuracy(%)
Tan et al. (2022) ResNet-18 (ANN) BiGRU (ANN) 72.1
Bulzomi et al. (2023) Modified Spiking ResNet + PLIF FC (Stateful Synapses) 60.2
Dampfhoffer et al. (2024) ResNet-18 (ANN) BiGRU (ANN) 75.1
Spiking ResNet-18 + PLIF FC (Stateful Synapses) 68.1
Spiking ResNet-18 + PLIF SpikGRU2+ (Bi-direction + Sigmoid Gates + Ternary Spikes) 75.3
Ours Modified Spiking ResNet-18 + Mul-free Channel-wise PSN FC (Stateful Synapses) 70.9
Table 3: Comparison with the state-of-the-art ANN and SNN methods on the DVS-Lip dataset.
5 Experiments
In this section, we evaluate the mul-free channel-wise PSN on various kinds of datasets. We conduct the ablation experiments on the order $k$ and demonstrate that the sawtooth dilations can compensate the long-term dependencies learning ability. Finally, we provide a training speed comparison to validate the efficiency of the autoselect algorithm.
5.1 Learning Long-Term Dependencies
We evaluate the long-term dependencies learning ability of mul-free channel-wise PSN in three widely used classification tasks, including the Spiking Heidelberg Digits (SHD) spoken digit dataset (Cramer et al., 2022), the sequential CIFAR dataset, and the high temporal resolution automatic lip-reading DVS-Lip dataset (Tan et al., 2022). These datasets cover the types of voices, images, and neuromorphic events.
Comparison between our methods and previous SOTA SNN methods on the SHD dataset are shown in Table 1. Specifically, we replace the LIF neurons in the SNN-delay architecture (Hammouamri et al., ) with our neurons. We achieve test accuracies of 95.31%, 95.62%, and 95.71% with sawtooth dilations and order $k=2,4,8$, respectively. To the best of our knowledge, these results represent the SOTA performance of SNN models on the SHD datasets.
Sequential image classification tasks have been commonly benchmarks to evaluate spiking neurons by (Yin et al., 2021; Fang et al., 2023b; Chen et al., 2024b). In these tasks, images are fed into the model column by column, and the number of time-step is equal to the width of the images. We also conduct experiments on sequential CIFAR10 and CIFAR100 datasets. To ensure fairness, we fully employ the network architecture and hyperparameters as (Fang et al., 2023b), only replacing the spiking neurons with ours. The results are shown in Table 2, where the data for PMSN is sourced from (Chen et al., 2024b), maintaining the same architecture as well, while the data for other neurons is sourced from (Fang et al., 2023b). On the sequential CIFAR10 dataset, our mul-free channel-wise PSN outperforms PSN by $2.72\%$ and PMSN by $0.2\%$. Additionally, on the sequential CIFAR100 dataset, it surpasses PSN by $4\%$ and PMSN by $0.13\%$. Notably, the order of our neurons here we report is 16, while the order of sliding PSN and masked PSN is 32, 2×2$\times$2 × than us.
Furthermore, we demonstrate the capability of mul-free channel-wise PSN in processing complex neuromorphic DVS-Lip dataset, which comprises 100 classes and consists of 19871 samples, each containing approximately $10^{4}$ events. These events are generated with a spatial resolution of $128\times 128$ pixels and a temporal resolution at the microsecond level. Half of the words in the dataset are visually similar pairs in the LRW dataset (Chung and Zisserman, 2017) (e.g., ”America” and ”American”). The training and testing sets are derived from different speakers, posing a challenge for the model to exhibit robust generalization capabilities with respect to speaker characteristics.
Currently, the SOTA accuracy of 75.3% on the DVS-Lip dataset is achieved by (Dampfhoffer et al., 2024) using a Spiking ResNet-18 with the channel-wise PLIF neurons fronted, a SpikGRU2+ backend, and events are integrated into 90 frames ($T=90$). In our experiments, we introduce several modifications to the frontend. We replace the PLIF neurons with our neurons and remove spiking neurons from the pooling layers, referring to this architecture as Modified Spiking ResNet-18. As Table 3 shows, our method achieves 70.9% accuracy and is only second to (Dampfhoffer et al., 2024) with SpikGRU2+ backend. It is worth noting that SpikGRU2+ is bi-directional with two groups of separate hidden states, employs sigmoid gates with floating activations, and outputs ternary spikes ($-1,1,0$), which is not a pure SNN module and might be difficult to be deployed to neuromorphic chips. The accuracy we report here is based on the neuron order $k=2$ and sawtooth dilations, indicating that our method can effectively capture rich historical information with a small order even when handling tasks involving long-time sequences.
5.2 Ablation Study
To validate that our neurons can effectively approximate the effect of a larger receptive field with a smaller order $k$ through sawtooth dilations, we conduct ablation experiments on the sequential CIFAR100 and pixel CIFAR10 classification tasks.
Figure 6 (a) illustrates the accuracy curves of mul-free channel-wise PSN and other neurons on the sequential CIFAR100 dataset, with the highest accuracy marked by a red ★. When the order is 2, the accuracy of our neuron already significantly surpasses the whole PSN family. Furthermore, when the order increases to 3 or more, the accuracy remains roughly around 66%. It is evident that our neuron is more robust than the sliding PSN, as it does not exhibit the issue of fluctuating accuracy while increasing order.
To evaluate the effectiveness of sawtooth dilations, we conduct an ablation study on the pixel CIFAR10 classification task. In this task, images are flattened into one-dimensional vectors as time series inputs to the network. Thus, the number of time-step is $T=1024$. We adopted the same network structure as (Chen et al., 2024b). Figure 6 (b) illustrates the accuracy curves with/without sawtooth dilations. It can be observed that the accuracy with sawtooth dilations is consistently higher than that without. Additionally, when the order $k$ is small, which is a practical case for deployment, the accuracy of our neuron with sawtooth dilations is much higher. These results validate that the sawtooth dilations compensate the effect of large receptive fields when using a small $k$.
Figure 6: The order-accuracy curves on (a) the sequential CIFAR100 and (b) the pixel CIFAR10.
Figure 7: Comparison of training speed on CIFAR100.
5.3 Training Acceleration
We compare the training speed of PSN and the mul-free channel-wise PSN implemented by the autoselect Algorithm 1. The naive manual implementations, using reshape operations and PyTorch’s 1-D convolutions for neuron layers and time batch dimension fusion or the vectorizing map for stateless layers, are also compared. The experiments are carried out on a Debian GNU/Linux 11(bullseye) server with an Intel(R) Xeon(R) Platinum 8336C CPU, a Nvidia A100-SXM4-80GB GPU and 32GB RAM. We set the batch size as 128.
The training duration ($s$/epoch) of different neurons under different order $k$ on CIFAR100 is shown in Figure 7. Note that both PSN and sliding PSN are implemented by matrix multiplication in GPUs (Fang et al., 2023b), their speeds are identical and decoupled with $k$. For the sake of easy reading, we plot the neurons with the same $k$ in similar styles. The results show that our autoselect algorithm greatly improves the efficiency of mul-free channel-wise PSN and achieves a much faster training speed than the naive manual implementations. When $T\leq 4$, our method is comparable to PSN/Sliding PSN. In this case, the matrixs are nearly strips in PSN/Sliding PSN because $T$ is much less than other dimensions, causing inefficient matrix multiplications. When $T$ continuously increases, our method is slower, which is caused by the fact that the quantization induces some additional overhead, and memory reading/writing caused by reshape or vectorizing map operations for processing inputs/outputs in our SNNs is slower than the dimension fusion. Nonetheless, the speed gaps are not significant. The high task accuracy and hardware-friendly advantages make mul-free channel-wise PSN a strong alternative for PSN and sliding PSN.
6 Conclusion
In this paper, we introduce a novel parallelizable spiking neuron model named mul-free channel-wise PSN, which employs the channel-wise convolutions to process the input sequences, avoids the large neuron order by sawtooth dilations, and gets rid of floating multiplications by efficient bit shift operations. The considerations of accelerating the training of SNNs with the proposed neuron models are also discussed in detail. Experimental results demonstrate that mul-free channel-wise PSN achieves significant performance improvements in temporal classification tasks, showcasing its superior capability to capture long-term dependencies. Our methods solve the dilemma of performance and computational costs of spiking neuron models, and our acceleration methods will benefit future research as a practical reference.
Appendix
A    Acceleration Details of Spiking Neuron Layer
A.1    Time-first/last + Vmap + MM
In Eq.(17), weight of the standard 1D convolution $\bm{W}_{q}$ is shaped as $[C,k]$. The standard 1D convolution operation could be implemented by matrix multiplication and vectorizing map. When the input sequence $\bm{X}\in\mathbb{R}^{T\times N}$ arrives, where the sequence length $T$ is known, for the time-first data layout, the weight matrix $\bm{A}\in\mathbb{R}^{C\times T\times T}$ can be generated as:
$\bm{A}[:][i][j]=\begin{cases}\begin{aligned} W_{q}[:][k-1-\frac{i-j}{d}],\ &i-d(k-1)\leq j\leq i\\ &\&\ (i-j)\%d=0\end{aligned}\\ 0,\text{otherwise}\end{cases},$ (19)
where [:]delimited-[]:[:][ : ] means the slice operation.
Similarly, for the time-last data layout, the weight matrix $A\in\mathbb{R}^{C\times T\times T}$ can be generated as:
$\bm{A}[:][j][i]=\begin{cases}\begin{aligned} W_{q}[:][k-1-\frac{i-j}{d}],\ &i-d(k-1)\leq j\leq i\\ &\&\ (i-j)\%d=0\end{aligned}\\ 0,\text{otherwise}\end{cases}.$ (20)
Applying the vectorizing map to the input sequence and the weight matrix across the channel dimension, the membrane potential $\bm{H}$ can be calculated through the matrix multiplication operation over channels in parallel:
$\bm{H}[c]=\begin{cases}\bm{A[c]X[c]},\text{time-first layout}\\ \bm{X[c]A[c]},\text{time-last layout}\end{cases}.$ (21)
.0.1 A.2    Time-first/last + Custom CUDA Kernel
Suppose $\bm{X}$ is the input sequence, $\bm{H}$ is the hidden states, and $\bm{\delta^{H}}$ is the gradient with respect to $\bm{H}$, obtained by automatic differentiation in PyTorch, all of them are shaped as $(T,N,C,H,W)$ or $(N,H,W,C,T)$. Suppose $\bm{W}$ and $\bm{b}$ are the weight and bias of the convolution, shaped as $[C,k]$ and $[C]$, respectively.
The process of forward propagation can be represented as:
$\bm{H}=\text{pad}(\bm{X},(k-1,0))\star\bm{W}+b,$ (22)
where pad represents padding $k-1$ zeros on the left of $\bm{X}$ over the time dimension $T$, and $\star$ denotes the convolution operation on the $T$ dimension of $X$ using $W$.
The process of backward propagation can be represented as:
$\frac{\partial L}{\partial\bm{X}}$ $=\text{pad}(\bm{\delta^{H}},(0,k-1))\star\text{flip}(\bm{W}),$ (23)
$\frac{\partial L}{\partial\bm{W}}$ $=\text{pad}(\bm{X},(k-1,0))\star\bm{\delta^{H}},$ (24)
$\frac{\partial L}{\partial b}$ $=\sum_{t,n,h,w}\bm{\delta^{H}}_{t,n,n,w},$ (25)
where $\text{flip}(\bm{W})$ represents flip $\bm{W}$ left and right along the $k$ dimension.
Beyond PyTorch (cuDNN), a custom CUDA implementation for two data layouts is also considered. To avoid the reshape and incident memory copying, we design CUDA kernels by OpenAI Triton for processing both data layouts directly. Specifically, we manually implement the Eqs.(22)-(25) using the Triton framework. We design a custom autograd function, where the aforementioned kernel functions are called in the forward and backward methods. Noting that the convolution operation in Eq.(23) is consistent with that in Eq.(22), so the triton kernel function remains the same. Taking the time-first layout as an example, Eq.(22) and (24) could be implemented as Algorithms 2 and 3, respectively.
Algorithm 2 Triton forward kernel
Input: The input sequence pointer $X_{ptr}$, weight matrix pointer $W_{ptr}$, the output sequence pointer $H_{ptr}$, point to the first address of tensors, shaped ad $[T+k-1,N,C,H,W]$, $[C,k]$ and $[T,N,C,H,W]$, respectively.
[…]
2:  Calculate the offset $\bm{X}_{offset}$, $\bm{W}_{offset}$ and $\bm{H}_{offset}$ of each thread, shaped as $[BN,BC,T,k]$, $[1,BC,1,k]$ and $[BN,BC,T]$, respectively
3:  Load values of $X_{ptr}+\bm{X}_{offset}$ and $W_{ptr}+\bm{W}_{offset}$ from memory to SRAM tensors $\bm{X}$ and $\bm{W}$
4:  Utilizing the broadcasting mechanism, perform the element-wise multiplication of $\bm{X}$ and $\bm{W}$
5:  Sum the output $\bm{H}$ along the $k$ dimension
6:  Store the values of $\bm{H}$ to $H_{ptr}+\bm{H}_{offset}$ address
Algorithm 3 Triton grad of weight kernel
Input: The grad of output pointer $O_{ptr}$, the input sequence pointer $X_{ptr}$, the grad of weight pointer $W_{ptr}$, point to the first address of tensors, shaped ad $[T,N,C,H,W]$, $[T+k-1,N,C,H,W]$ and $[C,k]$, respectively.
[…]
2:  Calculate the offset $\bm{O}_{offset}$, $\bm{X}_{offset}$ and $\bm{W}_{offset}$ of each thread, shaped as $[BN,BC,T,1]$, $[BN,BC,T,k]$ and $[BC,k]$, respectively
3:  Load values of $O_{ptr}+\bm{O}_{offset}$ and $X_{ptr}+\bm{X}_{offset}$ from memory to SRAM tensors $\bm{O}$ and $\bm{X}$
4:  Utilizing the broadcasting mechanism, perform the element-wise multiplication of $\bm{O}$ and $\bm{X}$
5:  Sum the grad of weight $\bm{W}$ along the $T$ and $k$ dimensions
6:  Atomic add the values of $\bm{W}$ to $W_{ptr}+\bm{W}_{offset}$ address
B    Acceleration Details of Stateless Layer
Dataset Optimizer Batch Size Epoch Learning Rate Scheduler
Sequential CIFAR10/100 AdamW 128 256 0.001 CosineAnnealingLR
Pixel CIFAR10 AdamW 128 128 0.001 CosineAnnealingLR
SHD Adam (wd=1e-5) 256 150 0.001 for weights 0.1 for delay CosineAnnealingLR for weights OneCycleLR for delay
DVS-Lip Adam (wd=1e-4) 32 200 fixed 3e-4 for 0-100 epochs (1e-4, 5e-6) for 100-200 epochs CosineAnnealingLR
Table 4: Training hyper-parameters for different datasets.
Stateless layers include the convolutional, batch normalization, pooling, and linear layers. When using the time-first data layout, the stateless layers can be accelerated by fusing the time dimension and the batch dimension in SpikingJelly. More specifically, the data layout changes as $(T,N,*)\rightleftharpoons(TN,*)$ before and after processing of the stateless layers. Then GPUs regard the time-step dimension as the batch dimension, leading to fully parallel computing over time-steps. It is worth noting that the dimension fusion is nearly no cost because the time and batch dimensions are physically adjacent in memory. The reshape operation only changes the view of tensors and does not involve the memory copying.
When using the time-last layout, the dimension fusion method of SpikingJelly cannot be applied expect for the batch normalization layer, which only requires that the channel dimension should be the 1-th dimension. Both layouts can be satisfied by reshape without additional memory R/W. For the convolutional and pooling layer, we introduce two new methods, the vectorizing map provided in PyTorch and the high-dimension convolution/pooling that has been used in the Lava framework, a software framework for neuromorphic computing.
The vectorizing map vectorizes the stateless layers to process the input sequence with the $(N,...,T)$ layout over the last dimension $T$, then the computation over time-steps is in parallel. This method actually implies a reshape operation $(N,...,T)\rightleftharpoons(T,N,...)$ when spliting and concatenating the sequence. The high-dimension convolution/pooling use the $(n+1)-$D convolution/pooling to implement the $n$-D convolution/pooling with weight as 1 and stride as 1 in the additional dimension, which is similar to using 2-D convolution to implement Eq.(17) discussed before. This method is also in parallel, while the main drawback is that the high-dimension convolution/pooling is complex and not as efficient as the dimension fusion method [Fang et al., 2023a].
C    Neuron Quantization
To alleviate the internal covariate shift along the temporal and batch dimension, increase the numerical stability of the model, we use the batch normalization to implement the learnable threshold $V_{th}$. Eq.(18) could be rewrite as:
$S[t][c]=\Theta\left(\gamma[c]\frac{H[t][c]-\mu_{\mathcal{B}}[c]}{\sqrt{\sigma^{2}_{\mathcal{B}}[c]+\epsilon}}+\beta[c]\right),$ (26)
where $\bm{\gamma}\in\mathbb{R}^{C}$ and $\bm{\beta}\in\mathbb{R}^{C}$ are the learnable weight, initialized as $1$ and $-1$. $\bm{\mu}_{\mathcal{B}}\in\mathbb{R}^{C}$ and $\bm{\sigma}^{2}_{\mathcal{B}}\in\mathbb{R}^{C}$ are the mean and variance of the input over the dimension $C$. Specifically, at train time, they are the mean and biased variance of the input sequence; at inference time, they are the moving average of the mean and unbiased variance of the input sequence on the training stage, which means $\bm{\mu}_{\mathcal{B}}$ and $\bm{\sigma}^{2}_{\mathcal{B}}$ are invariant during inference.
Since our quantization goal is to use the efficient bit shifting operation to replace the multiplication, the convolution layer and the batch normalization layer could be fused to reduce computation during inference, so we need to quantize the fused weights during training. The formula for the fusing of convolution and batch normalization can be represented as follows:
$\bm{W}_{f}$ $=\frac{\bm{\gamma}}{\sqrt{\bm{\sigma}^{2}_{\mathcal{B}}+\epsilon}}\cdot\bm{W},$ (27)
$\bm{b}_{f}$ $=\bm{\beta}-\frac{\bm{\gamma}\cdot\bm{\mu}_{\mathcal{B}}}{\sqrt{\bm{\sigma}^{2}_{\mathcal{B}}+\epsilon}}.$ (28)
Thus, $\bm{W}$ in Eq.(13) is actually $\bm{W}_{f}$ in Eq.(27). To implement the quantization of fused weight, during the training stage, input sequence $\bm{X}$ is first passed to the convolutional layer, resulting in the intermediate variable $\bm{T}$ to update the $\bm{\mu}_{\mathcal{B}}$ and $\bm{\sigma}^{2}_{\mathcal{B}}$ in Eq.(27) and Eq.(28). Then, we use $\bm{W}_{f}$ as $\bm{W}$ in Eq.(13) and $\bm{b}_{f}$ as $\bm{V}_{t}h$ in Eq.(18), perform the convolution operation on the input $\bm{X}$ twice. After the training is completed, $\bm{\mu}_{\mathcal{B}}$ and $\bm{\sigma}^{2}_{\mathcal{B}}$ is fixed, so we could directly use the quantized $\bm{W}_{f}$ and $\bm{b}_{f}$ as the weight and bias of the convolution layer, performing the convolution operation only once.
D    Network Structure
Dataset Network structure
Sequential CIFAR10/ CIFAR100 {{c128k3s1-BN-SN}*3-APk2s2}*2-FC256-SN-FC10/100
Pixel CIFAR10 FC128-BN-SN-{RB128}*2-APk4s4-FC256-BN-SN-{RB256}*2-FC10
SHD {Dcls256-BN-SN-DP}*2-Dcls20-LIF(Vth=1e9)
DVS-Lip 3D_c64k577s122p233-APk3s2p1-{RB64}*2-{RB128}*2-{RB256}*2-{RB512}*2-AAPk1-DP-Stateful FC 100
Table 5: Network structure for different datasets.
Table 5 illustrates the details of the network structure for different datasets. c128k3s1 represents convolution layer with output channels $=128$, kernel size $=3$ and stride $=1$, BN is the batch normalization. SN is the mul-free channel-wise PSN, APk2s2 is the avg-pooling layer with kernel size $=2$ and stride $=2$, FC256 represents the fully connected layer with output feature $=256$. RB128 is the residual block with output channels $=128$, Dcls256 is the dilated convolution with learnable spacings with output channels $=256$, DP is the dropout layer. LIF(Vth=1e9) represents an LIF spiking neuron the threshold $=1e9$, and the membrane potential is the output, which could be thought of the moving average of the input current. 3D_c64k577s122p233 represents the 3D convolution layer with output channels $=64$, kernel size $=(5,7,7)$, stride $=(1,2,2)$ and padding $=(2,3,3)$. AAPk1 is the adaptive avg-pooling layer with output feature $=1$. Stateful FC 100 is a FC layer with stateful synapses.
E    Setting of Experiments
The main hyper-parameters for different datasets are shown in Table 4. Other training options are listed as follows.
Sequential CIFAR The data augmentation techniques include random mixup with $p=1$ and $\alpha=0.2$, random cutmix with $p=1$ and $\alpha=1$, random choice between the two mix methods with $p=0.5$, random horizontal flip with $p=0.5$, trivial augmentation, normalization, random erasing with $p=0.1$, and label smoothing with the amount $0.1$ [Fang et al., 2023b]. The number of channels is $128$. The surrogate function is the arctan surrogate function $\sigma(x)=\frac{\alpha}{2(1+(\frac{\pi}{2}\alpha x)^{2})}$ with $\alpha=2$.
Pixel CIFAR All is the same as Sequential CIFAR.
SHD No specific augmentation method is implemented. The surrogate function is also the arctan surrogate function with $\alpha=5$.
DVS-Lip The data augmentation techniques include center cropped size $=96\times 96$, then random cropped size $=88\times 88$, random horizontal flip with $p=0.5$, 2D spatial mask with mask num $=4$ and maximum length$=20$, random choice between zoom in and zoom out with $p=0.5$ and max scale $=26$, temporal mask with mask num $=6$ and maximum length $=18$ [Dampfhoffer et al., 2024]. The surrogate function is $\sigma(x)=\frac{1}{1+\alpha x^{2}}$ with $\alpha=10$.
F    Experimental Results
The original data for Figure 6 and Figure 7 are shown in Table 6 and Table 7, respectively. Notably, the training duration of our neuron is with the quantization operation, i.e. perform the convolution operation twice, so it is inherently slower than PSN on the training stage.
Order Dataset sequential CIFAR100 pixel CIFAR 10 (w/o dilation) pixel CIFAR 10 (w/ dilation)
1 50.24 45.15 45.33
2 63.25 66.21 68.96
3 65.21 69.33 74.21
4 65.77 72.97 76.82
5 66.45 73.90 77.31
6 65.96 75.44 78.92
7 66.58 75.21 79.49
8 66.97 76.72 81.71
9 66.48 79.97 82.67
10 66.38 79.75 82.36
11 66.87 79.80 82.74
12 66.53 79.39 83.24
13 66.06 79.65 82.85
14 67.15 80.07 83.58
15 66.60 79.37 83.32
16 66.21 79.83 83.65
17 66.73 80.13 84.89
18 66.32 80.51 84.84
19 66.41 80.89 84.28
20 66.10 80.56 84.05
21 66.41 81.19 84.04
22 66.23 81.44 84.43
23 66.45 81.00 84.01
24 65.98 80.71 84.11
25 66.56 80.78 83.49
26 66.53 81.26 82.85
27 66.30 81.61 82.60
28 66.40 81.58 82.33
29 66.59 81.70 82.81
30 66.40 80.90 82.71
31 66.62 82.11 82.59
32 66.36 81.84 83.07
Table 6: Original test accuracy (%) of Figure 6.
Method T 2 4 8 16 32
PSN/SlidingPSN 17.86 23.04 27.47 44.18 88.12
Time_first(k=2) 28.23 42.10 493.01 623.1 881.87
Time_first(k=4) 43.63 771.39 888.15 1154.12
Time_first(k=8) 1379.37 1413.37 1696.01
Time_last(k=2) 30.36 42.43 286.33 369.94 547.64
Time_last(k=4) 52.44 493.25 622.84 845.38
Time_last(k=8) 1379.70 1422.11 1657.56
Autoselect(k=2) 15.06 25.33 45.82 84.91 181.06
Autoselect(k=4) 24.67 45.59 77.60 231.65
Autoselect(k=8) 46.40 78.22 193.85
Table 7: Original training duration (s/epoch) of Figure 7.
In Figure 4, we mention that the gradient of Eq.(15) is jump points, which is detrimental to the network. On the sequence CIFAR dataset, we find that the network is still able to learn quite well. However, on the DVS-Lip dataset, as shown in Figure 8, using the original ste gradient causes the network to crash, resulting in the training and testing accuracy suddenly dropping to $1\%$. The reason is that the gradients appear to be the Not a Number (NaN) values.
Figure 8: The training and testing accuracy curves with gradient of Eq.(14) on the DVS-Lip Dataset.